what is the answer of 4x+y=6 and y=2x
Answer by Cromlix(4381)

4x+y=6 and y=2x
Substitute y = 2x into
4x + y = 6
4x + 2x = 6
6x = 6
x = 1
y = 2x
y = 2*1
y = 2
{1,2}
Hope this helps.
:-)